### v2

@@ 4.1 Learning-based interest rate controller
When each user (borrower or lender) is infinitesimal, their individual supply or demand has no impact on the resulting interest rate $r_{t}$. Consequently, their bidding strategy does not influence the price they pay or receive and they are incentivized to report their true valuation. However, in general cases, the situation changes. For example, when lenders are inelastic and the supply is fixed, with only borrowers participating, the auction reduces to a multi-unit demand auction with a uniform price, which is known to be non-truthful and susceptible to demand reduction [1, 21]. The VCG (Vickrey-Clarke-Groves) auction is known to be truthful in this context; however, it assigns different prices to different participants, which is not feasible in a peer-to-pool-to-peer setting. However, a peer-to-peer protocol can be attached on top of a pool, subject to constraint on the peer-to-peer contracting (e.g. Morpho protocol [11]). But whether a VCG-like truthful mechanism can be implemented using Morpho is an open question and beyond the scope of our work.
Theorem 4.1
@@
4.2 Static Interest Rate Curves
Unlike adaptive algorithms, which are vulnerable during the learning phase, static curves are memoryless, reducing the scope for adversarial manipulation. However, strategic users can still exploit their knowledge of the fixed curve’s structure. We identify a tactic called strategic withholding, where lenders supply less than their full capacity to raise the utilization rate, thus increasing the pool’s interest rate and boosting their returns. Similarly, borrowers might borrow less than their maximum capacity to lower their overall interest costs by fulfilling the remaining demand externally. This strategy was first introduced by [33].
@@ 4.2 Static Interest Rate Curves
Let a strategic borrower $\mathcal{A}^{b}$, with external rate $r^{b}$, control $\delta_{b}$ of the total demand $B$, and a strategic lender $\mathcal{A}^{l}$, with external rate $r^{l}$, control $\delta_{l}$ of the total supply $L$. Both aim to maximize their utility by adjusting how much they participate in the pool. The full utility functions for strategic borrowers and lenders are provided in Appendix [0.A.6].
@@
Theorem 4.2
Given the static curve described in 8, with fixed truthful demand $(1-\delta_{b})B$ and fixed truthful supply $(1-\delta_{l})L$, and the remaining demand $\delta_{b}B$ and supply $\delta_{l}L$ controlled by strategic borrowers and lenders, the Adversarial Impact is bounded by:
$AI^{\mathcal{P}}(\delta_{b},\delta_{l})\leq\max\left\{\frac{B\delta_{b}R_{\text{slope2}}}{L(1-U^{*})},\frac{B\delta_{l}R_{\text{slope2}}}{L(1-\delta_{l})(1-U^{*})}\right\}$ (14)
@@
5 Evaluation
@@
Token DAI USDC WETH
Average Error $7.3\%$ $3.5\%$ $4.4\%$
@@
Table 1: 10-Step ahead normalized error of Aave demand and supply estimates
@@
Empirical validation of demand and supply models: We applied our RLS-based algorithm to demand and supply data from Aave market pools, analyzing the relationship between interest rates and the demand and supply curves in 3-hour intervals. Specifically, we fitted the data (Jan-Jul 2023, RLS with $\rho=0.8$) to our demand and supply models (Equations 10 and 9) and estimated the parameters $\mathtt{a}_{t}^{b}$, $\mathtt{b}_{t}^{b}$, $\mathtt{a}_{t}^{l}$, and $\mathtt{b}_{t}^{l}$ over time. To evaluate the accuracy of these parameter estimates, we predicted demand and supply for the next 10 intervals and summarized the prediction errors in Table 1. The relatively low error across the three main pools demonstrates the effectiveness of our model and methodology. Visualizations of parameter evolution are provided in the Appendix.0.B.1.
(a) Comparison of utilization between RLS-based controller and Aave’s static curves
@@
(b) Impact of transition noise on average utilization error for RLS and Aave controllers
Figure 2: Comparing our interest rate controller with Aave on market conditions learnt from on-chain data
Utilization Error We evaluated the performance of our RLS-based controller in regulating utilization compared to Aave’s static curves. Using demand and supply parameters estimated from Aave data, we replicated real user behavior and set interest rates with our RLS algorithm. Figure [2(a)] shows that our controller maintains utilization close to the target, even with changing demand and supply. We also tested robustness by simulating demand and supply curves with their parameters performing a Gaussian random walk with a specific transition noise. As shown in Figure [2(b)], the RLS controller consistently outperforms Aave’s, even under high transition noise levels.
Forgetting factor and estimation error The forgetting factor governs the trade-off between adaptivity and precision; therefore, depending on the dynamics of the demand and supply parameters, different forgetting factors should be used. Figure 6 in Appendix [0.B.2] illustrates the mean squared error (MSE) of the parameters $\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$ estimated by the RLS-based algorithm. The parameters were generated using a random walk with Gaussian noise, with a frequency of change every $T=50$ steps. In this scenario, the optimal forgetting factor, which minimizes the overall MSE across all parameters, is approximately $\rho=0.82$.
Figure 3: ETH volatility over time, with the controller dynamically adjusting the collateral factor. The bottom plot compares the actual liquidations resulting from this adjusted collateral factor with the target expected liquidations, $LT_{t}=0.9$.
@@
Controlling liquidation Figure 3 shows our risk controller adjusting the collateral factor to liquidation threshold ratio as Ethereum’s volatility changes. The goal is to maintain an average liquidation rate of 1% (orange) or 0.1% (blue), with a fixed liquidation threshold of $LT_{t}=0.9$. The controller lowers the collateral factor during high volatility to prompt borrowers to add collateral and avoid liquidation.
6 Conclusion
@@
Adaptivity and Robustness In this work, we introduced a model of non-stationary borrower and lender behavior that aligns with empirical data, along with an interest rate controller designed to maintain stability under dynamic market conditions. We also characterized the protocol’s responsiveness to market changes and analyzed the limits of this adaptivity when facing adversarial manipulation.
Limitations and future work Our model has three key limitations, leaving room for future research. First, we assume that lending does not affect the collateral’s price, which may not hold true for low-liquidity assets. Second, we focused on a single pair of lent and collateral assets. A more comprehensive model would adjust interest rates for a broader range of assets, considering interactions between multiple borrowed and collateralized assets, with potentially correlated price and volatility movements. Thirdly, Section 4 demonstrates the vulnerability of a fully automated adaptive approach to adversaries, calling for manual/governance guardrails while bringing automation to DeFi.
Appendix 0.A Supplementary explanations and results
0.A.1 Details of Market Participants
Lenders Lenders deposit asset $\mathcal{A}_{l}$ to earn interest on the amounts borrowed by the borrowers. In this analysis, we focus on a lending protocol where all accrued interest goes directly to the lenders. Unlike many existing platforms, which reserve a portion of the interest for a fund to cover potential defaults and distribute some as rewards to token holders [3, 6], the platform we study has no such reserves. As a result, any borrower default directly reduces the deposits of lenders, transferring the risk of default from the protocol to the lenders. Let $L_{t}(i)$ be the deposit of lender $i$ in $\mathcal{A}_{l}$, and let $L_{t}$ be the total deposits. The utilization rate, $U_{t}=\frac{B_{t}}{L_{t}}$, expresses the fraction of deposits currently borrowed. Lenders earn interest at rate $r_{t}$ on the portion of their deposits that are utilized. In the event of a default, each lender’s loss is proportional to their original deposit relative to the total pool.
0.A.2 Utility Function of Strategic Users
Here we outline the incentives of strategic users and their utility function. At any time $t$, a strategic lender $i$ allocates a portion (or potentially all) of their supply $\hat{L}_{t}(i)\leq L_{t}(i)$ to $\mathcal{P}$ and allocate the remaining supply, $L_{t}(i)-\hat{L}_{t}(i)$, to an alternative external market that offers a rate $r^{l}_{t}(i)$. The lender’s objective is to maximize cumulative utility over time, derived from expected interest rates from both $\mathcal{P}$ and the alternative market. Formally, the cumulative utility over time for the lender is expressed as:
$\mathbb{E}\left[\sum_{t=0}^{\infty}\left(\hat{L}_{t}(i)\,r_{t}\,U_{t}+\left(L_{t}(i)-\hat{L}_{t}(i)\right)\,r^{l}_{t}(i)\right)\right]$ (15)
$\mathbb{E}\left[\sum_{t=0}^{\infty}\left(\hat{B}_{t}(i)\,r_{t}+\left(B_{t}(i)-\hat{B}_{t}(i)\right)\,r^{b}_{t}(i)\right)\right]$ (16)
0.A.3 Alternate objective for demand/supply balancing
@@
$r^{*}_{\text{rev}}\coloneqq\arg\max_{r}\left[f(r;\theta)+g(rU;\omega)\right]\quad\text{subject to}\;U<U_{\text{max}}\;\text{where}\;U=\frac{f(r;\theta)}{g(rU;\omega)}$ (17)
@@
0.A.4 Optimizing interest rate to maximize revenue
@@
Theorem 0.A.1
Given the user behavior models described in Equations 10 and 9, the revenue maximization problem, as described in Equation 17, can be formulated as follows:
$\max_{r}\;\big{(}-\mathtt{a}^{b}\,r+\mathtt{a}^{l}\,r\,U\big{)}$ (18)
@@
$r_{\text{rev}}^{*}=\begin{cases}\frac{\mathtt{b}^{b}+\mathtt{b}^{l}U_{\text{max}}}{\mathtt{a}^{b}+\mathtt{a}^{l}(U_{\text{max}})^{2}}&\text{if}\frac{\mathtt{a}^{b}\,\mathtt{b}^{l}(1+\sqrt{1-4\,\mathtt{a}^{b}\,\mathtt{a}^{l}})}{\mathtt{a}^{l}\left(\mathtt{b}^{b}+\sqrt{(\mathtt{b}^{b})^{2}+4\,(\mathtt{a}^{b}\,\mathtt{b}^{l})^{2}}\right)}\geq U_{\text{max}}\\ \frac{\mathtt{b}^{b}+\sqrt{(\mathtt{b}^{b})^{2}+4\,(\mathtt{a}^{b}\,\mathtt{b}^{l})^{2}}}{2\,\mathtt{a}^{b}}&\text{O.W. }\end{cases}$ (19)
Informally speaking, replacing the expected optimal rate at line 5 of Algorithm 2 with Equation 19 will result in an RLS-based revenue-maximizing rate controller, offering similar convergence guarantees as those outlined in Theorem 3.1.
0.A.5 Formal abstraction of the learning-based controller

### v3

@@ 5.1. Learning-based interest rate controller
When each user (borrower or lender) is infinitesimal, their individual supply or demand has no impact on the resulting interest rate $r_{t}$. Consequently, their bidding strategy does not influence the price they pay or receive and they are incentivized to report their true valuation. However, in general cases, the situation changes. For example, when lenders are inelastic and the supply is fixed, with only borrowers participating, the auction reduces to a multi-unit demand auction with a uniform price, which is known to be non-truthful and susceptible to demand reduction (eco, [n. d.]; Engelbrecht-Wiggans and Kahn, 1998). The VCG (Vickrey-Clarke-Groves) auction is known to be truthful in this context; however, it assigns different prices to different participants, which is not feasible in a peer-to-pool-to-peer setting. However, a peer-to-peer protocol can be attached on top of a pool, subject to constraint on the peer-to-peer contracting (e.g. Morpho protocol (morpho, [n. d.])). But whether a VCG-like truthful mechanism can be implemented using Morpho is an open question and beyond the scope of our work.
Theorem 5.2.
@@
5.2. Static interest rate curves
Unlike adaptive algorithms, which are vulnerable during the learning phase, static curves are memoryless, reducing the scope for adversarial manipulation. However, strategic users can still exploit their knowledge of the fixed curve’s structure. We identify a tactic called strategic withholding, where lenders supply less than their full capacity to raise the utilization rate, thus increasing the pool’s interest rate and boosting their returns. Similarly, borrowers might borrow less than their maximum capacity to lower their overall interest costs by fulfilling the remaining demand externally. This strategy was first introduced by (Yaish et al., 2023).
@@ 5.2. Static interest rate curves
Let a strategic borrower $\mathcal{A}^{b}$, with external rate $r^{b}$, control $\delta_{b}$ of the total demand $B$, and a strategic lender $\mathcal{A}^{l}$, with external rate $r^{l}$, control $\delta_{l}$ of the total supply $L$. Both aim to maximize their utility by adjusting how much they participate in the pool. The full utility functions for strategic borrowers and lenders are provided in Appendix [A.4].
@@
Theorem 5.3.
Given the static curve described in 10, with fixed truthful demand $(1-\delta_{b})B$ and fixed truthful supply $(1-\delta_{l})L$, and the remaining demand $\delta_{b}B$ and supply $\delta_{l}L$ controlled by strategic borrowers and lenders, the Adversarial Impact is bounded by:
(16) $AI^{\mathcal{P}}(\delta_{b},\delta_{l})\leq\max\left\{\frac{B\delta_{b}R_{\text{slope2}}}{L(1-U^{*})},\frac{B\delta_{l}R_{\text{slope2}}}{L(1-\delta_{l})(1-U^{*})}\right\}$
@@
5.3. Mitigating the effects of manipulation
5.3.1. Ecosystem modification
In Theorem 5.2, we have shown that the interest rate of a lending pool that follows an adaptive algorithm is susceptible to adversarial manipulation via strategic withholding. We now see if there are ways of mitigating this kind of a manipulation.
Note that the core of the problem stems from the fact that the demand (or supply) of liquidity to the pool reacts in a delayed fashion to deviation of the demand (or supply) from the norm. By the “norm”, we mean the demand (or supply) that borrowers (or lenders) have for the asset if everyone was honest. This delay is captured by the parameter of demand (or supply) elasticity. A more elastic demand implies a more reactive market. This, in turn, implies a higher chance of the demand seen by the lending pool to be a true representation of market demand. The question of mitigating adversarial effects on demand and supply boils down to the question of increasing supply and demand elasticities of the underlying asset with respect to the interest rates. This increase can be achieved by the following strategies.
Naively, the elasticity of a pool can be boosted by exposing the pool to as much alternative trading volume as possible. The presence of multiple competing lending pools would increase elasticity, but is rarely in the hands of the lending pool designer. Additionally, the designer can replicate the pool across different blockchain platforms including Layer-1’s and Layer-2’s. The presence of fast and robust bridges across all of these platforms would help unify liquidity and encourage arbitrage and hence increase pool elasticity.
Alternatively, the presence of a deep derivatives market for interest rates of specific pools in a lending protocol will also help increase demand/supply elasticity. A derivatives market enables speculators and arbitrageurs to take note of discrepancies in what the futures markets expect the interest rates to be and what it actually is due to adversarial manipulation. In traditional finance, for example, the use of credit derivatives has been observed to be associated with an increased supply of bank credit (Hirtle, 2008). On the demand side, the use of credit derivatives such as interest rate swaps allow more sophisticated users to face a variable interest rate while less sophisticated users can fix their desired interest rate. This allows a user more capable of managing an adaptive interest rate to take on the risk entailed by such a rate. This user would also more sensitive in assessing whether the current interest rate offered by a pool is reasonable, and takes an action to borrow more if the rate is too low. This, in turn, would increase the demand elasticity of the interest rate market. Markets offering yield tokenization derivatives such as Pendle (Finance, 2022), or fixed rate lending/borrowing such as IPOR (Camas et al., 2021) effectively allow users in DeFi to enter interest rate swaps. Other protocols such as Timeswap (Ngo, 2022) effectively offer a futures market on interest rates. In principle, a sophisticated trader should be able to use the prices of swaps on these external markets and assess whether the adaptive interest rate offered on a lending pool is fair. A transparent valuation model for all such protocols is an interesting open problem and would help boost elasticity and hence protocol robustness for all lending pools, but is beyond the scope of this paper. We leave the pricing of such derivatives to future work.
5.3.2. Algorithmic mitigation
To address adversarial manipulation of the demand and supply functions, we propose a robust version of the RLS algorithm where the adaptive controller evaluates the plausibility of incoming data by comparing it to prior estimations of system dynamics. Data points that deviate significantly from expected patterns are given a lower weights in estimating the demand and supply functions. In particular, we employ the outlier-robust recursive least squares (RLS) estimator with a forgetting factor, as introduced in (Zou et al., 2000; Kovačević et al., 2016). This robust estimator minimizes the following modified loss function, an alternative to the standard RLS loss function with a forgetting factor in Equation 13:
(17) $J_{t}(\bm{\theta})=\sum_{\tau=0}^{t}\rho^{t-\tau}\phi\left((y_{\tau}-\mathbf{x}_{\tau}^{T}\bm{\theta})^{2}\right),$
where $\phi(.)$ is the M-estimate function. The M-estimate function is a redescending function, meaning its derivative decreases to zero as the residual $e$ becomes very large. Refer to Appendix A.6 for the description of the M-estimate function we adopt. Zou et al. show that minimizing the robust loss function in Equation 17 results in a recursive algorithm similar to the standard RLS with forgetting factor, except that the gain is updated as: $\mathbf{K}_{t}=\frac{q(e_{t})\mathbf{P}_{t-1}\mathbf{x}_{t}}{\rho+q(e_{t})\mathbf{x}_{t}^{T}\mathbf{P}_{t-1}\mathbf{x}_{t}},$ where $e_{t}\coloneqq(y_{t}-\mathbf{x}_{t}^{T}\hat{\bm{\theta}}_{t-1})^{2}$ and $q(e)\coloneqq\frac{d\phi(e)}{de}/e$.
6. Evaluation
@@
Token DAI USDC USDT
Average Error $2.1\%$ $1.3\%$ $1.2\%$
@@
Table 1. Normalized error of Compound demand and supply estimates from Feb 2024 - Feb 2025.
6.1. Empirical validation of demand and supply models:
@@
We applied our RLS-based algorithm to demand and supply data from Compound market pools, analyzing the relationship between interest rates and the demand and supply curves in 3-hour intervals. Specifically, we fitted the data (Feb 2024 - Feb 2025, RLS with $\rho=0.95$) to our demand and supply models (Equations 12 and 11) and estimated the parameters $\mathtt{a}_{t}^{b}$, $\mathtt{b}_{t}^{b}$, $\mathtt{a}_{t}^{l}$, and $\mathtt{b}_{t}^{l}$ over time. To evaluate the accuracy of these parameter estimates, we predicted demand and supply for the next timeslot and summarized the prediction errors in Table 1. The relatively low error across the three main pools demonstrates the effectiveness of our model and methodology. More details about our experiment are provided in the Appendix B.1.
(a) Robust RLS achieves near-zero rate deviation even at extreme ($\sigma_{\text{attack}}$).
@@
(b) Impact of transition noise on average utilization error for RLS and Aave controllers.
6.2. Interest rate controller
To evaluate our interest rate controller, we analyze two demand-supply models: one based on real-world Compound data (Section 6.1) and another using synthetic parameter trajectories. The latter follows a Gaussian random walk, where each parameter ($\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$) evolves every $T$ steps as $\mathtt{a}^{b}_{t+T}=\mathtt{a}_{t}^{b}+\epsilon_{t}$, with $\epsilon_{t}\sim\mathcal{N}(0,\sigma_{\text{trns}}\cdot|\mathtt{a}_{t}^{b}|)$. Demand and supply are then generated using Equations 11 and 12, incorporating normally distributed noise ($\sigma_{\text{noise}}$).
Pool Utilization MSE RLS Compound
DAI 0.0085 0.0141
USDC 0.0009 0.0079
USDT 0.0001 0.0032
Table 2. Utilization MSE comparison for RLS-based controller and Compound Static Curve.
$\sigma_{\text{trns}}$ Adaptive FF $\rho=0.85$ $\rho=0.9$ $\rho=0.99$
0.1 51.15 54.13 52.72 51.48
0.2 51.58 54.58 53.08 56.32
0.4 56.62 58.12 58.59 107.60
Table 3. RMSE of estimated demand function by RLS for different transition noises and forgetting factors.
Utilization error We evaluate our RLS-based controller against Compound’s static interest rate curves tuned for optimal utilization of $0.8$. Using Compound’s data (Feb 2024–Feb 2025), we infer $\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$ to simulate user behavior. Our controller archives significantly lower utilization mean squared error (MSE) across all pools (Table 3). We use $\rho=0.85$. To assess robustness, we test synthetic data. As $\sigma_{\text{trns}}$ increases, inducing greater demand-supply fluctuations, our controller incurs more errors but consistently outperforms the Aave/Compound static approach (Figure 2(b)). For details, see Appendix B.2.
Variable Forgetting Factor This experiment evaluates the impact of different forgetting factors on demand estimation under varying parameter evolution rates ($\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b}$) and examines whether the adaptive forgetting factor outperforms RLS with a fixed forgetting factor. We consider transition noise levels $\sigma_{\text{trns}}=0.1,0.2,$ and $0.4$. As shown in Table 3, the optimal fixed forgetting factor decreases with increasing noise: $\rho=0.99$ performs best at $\sigma_{\text{trns}}=0.1$, while $\rho=0.9$ and $\rho=0.85$ yield lower errors at $\sigma_{\text{trns}}=0.2$ and $0.4$, respectively. However, the adaptive forgetting factor consistently achieves the lowest error across all cases.
Robust RLS We assess the robustness of robust RLS against adversarial demand and supply manipulation. The metric of interest is the normalized rate deviation, measuring deviation from the optimal rate for $U^{*}=0.7$. Two attack strategies are considered. The first is an intermittent adversary, which randomly activates with probability 0.1, increasing the demand and supply noise variance up to $\sigma_{\text{attack}}$ (20%–300% of demand or supply). As shown in Figure 2(a), plain RLS exhibits sharp deviation growth, whereas robust RLS maintains near-zero deviation. The second strategy simulates a persistent adversarial borrower who inflates $\mathtt{a}_{t}^{b}$ by $\gamma_{\text{adv}}$ to exaggerate demand sensitivity, tricking the protocol into setting lower rates. With $T_{\text{attack}}=100$ and $\gamma_{\text{adv}}\in\{2,5,10,20\}$, plain RLS suffers a normalized rate deviation of up to 3.13, meaning the rate reaches 3.13 times its optimal value. Robust RLS keeps it below 0.5, demonstrating resilience.
6.3. Risk controller
@@
Figure 5 shows our risk controller adjusting the collateral factor to liquidation threshold ratio as Ethereum’s volatility changes. The goal is to maintain an average liquidation rate of 1% (orange) or 0.1% (blue), with a fixed liquidation threshold of $LT_{t}=0.9$. The controller lowers the collateral factor during high volatility to prompt borrowers to add collateral and avoid liquidation.
7. Conclusion
@@
Adaptivity and robustness In this work, we introduced a model of non-stationary borrower and lender behavior that aligns with empirical data, along with an interest rate controller designed to maintain stability under dynamic market conditions. We also characterized the protocol’s responsiveness to market changes and analyzed the limits of this adaptivity when facing adversarial manipulation. We suggested solutions to reduce adversarial manipulation.
Limitations and future work Our model has three key limitations, suggesting directions for future work. First, it assumes lending does not impact collateral prices, which may not hold for low-liquidity assets. Second, it considers only a single lent-collateral pair, whereas a broader model would account for multiple assets with correlated price and volatility movements. Third, we focus on Compound-style pools, where collateral remains idle, unlike Aave’s cross-asset model, which generates yield but adds complexity to the risk analysis of the protocol. Due to these risk management challenges, Compound opts for a simpler design, which we also adopt in our analysis (Guides, 2025). Finally, leveraging neural networks and deep reinforcement learning could enhance user modeling and protocol design. However, their computational inefficiency hinders on-chain use, even with rollups, and their lack of interpretability challenges financial transparency. Developing efficient, interpretable algorithms remains an open problem.
Appendix A Supplementary explanations about the model and the results
A.1. Alternate objective for demand/supply balancing
@@
(18) $r^{*}_{\text{rev}}\coloneqq\arg\max_{r}\left[f(r;\theta)+g(rU;\omega)\right]\quad\text{subject to}\;U<U_{\text{max}}\;\text{where}\;U=\frac{f(r;\theta)}{g(rU;\omega)}$
@@
A.2. Optimizing interest rate to maximize revenue
@@
Theorem A.1.
Given the user behavior models described in Equations 12 and 11, the revenue maximization problem, as described in Equation 18, can be formulated as follows:
(19) $\max_{r}\;\big{(}-\mathtt{a}^{b}\,r+\mathtt{a}^{l}\,r\,U\big{)}$
@@
(20) $r_{\text{rev}}^{*}=\begin{cases}\frac{\mathtt{b}^{b}+\mathtt{b}^{l}U_{\text{max}}}{\mathtt{a}^{b}+\mathtt{a}^{l}(U_{\text{max}})^{2}}&\text{if}\frac{\mathtt{a}^{b}\,\mathtt{b}^{l}(1+\sqrt{1-4\,\mathtt{a}^{b}\,\mathtt{a}^{l}})}{\mathtt{a}^{l}\left(\mathtt{b}^{b}+\sqrt{(\mathtt{b}^{b})^{2}+4\,(\mathtt{a}^{b}\,\mathtt{b}^{l})^{2}}\right)}\geq U_{\text{max}}\\ \frac{\mathtt{b}^{b}+\sqrt{(\mathtt{b}^{b})^{2}+4\,(\mathtt{a}^{b}\,\mathtt{b}^{l})^{2}}}{2\,\mathtt{a}^{b}}&\text{O.W. }\end{cases}$
Informally speaking, replacing the expected optimal rate at line 4 of Algorithm 2 with Equation 20 will result in an RLS-based revenue-maximizing rate controller, offering similar convergence guarantees as those outlined in Theorem 4.1.
A.3. Formal abstraction of the learning-based controller